OpenStudy (anonymous):

prove that (sin 2theta +1)/cos2theta =(cos theta +sin theta)/(cos theta -sin theta)

OpenStudy (turingtest):

is that \[\sin(2\theta)\]or\[\sin^2\theta\]?

OpenStudy (anonymous):

sin(2θ)

OpenStudy (anonymous):

Multiply the numerator and denominator of the right hand side by \( (\cos \theta + \sin \theta )\) and the result should follow instantly.

OpenStudy (anonymous):

alright lemme try

OpenStudy (anonymous):

\[[( \sin 2\theta +1)/ \cos2\theta ] \times(\cos \theta +\sin \theta)\]

OpenStudy (anonymous):

is this what you asked me to do?

OpenStudy (turingtest):

no that's the left-hand side do the other side instead

OpenStudy (anonymous):

OHH I SEEE

OpenStudy (anonymous):

i got the answer hurray!

OpenStudy (anonymous):

i have another question i need help with

OpenStudy (anonymous):

prove that \[\cos 4x =8\cos ^{4}x - 8\cos ^{2}x +1\]

OpenStudy (anonymous):

New thread please :)

OpenStudy (anonymous):

ok